### v3

@@ Abstract
Mining Software Repositories (MSR) has become an essential activity in software development. Mining architectural information (e.g., architectural models and views) to support architecting activities, such as architecture recovery and understanding, has received significant attention in recent years. However, there is a lack of clarity on what literature on mining architectural information is available. Consequently, this may create difficulty for practitioners to understand and adopt the state-of-the-art research results, such as what approaches should be adopted to mine what architectural information in order to support architecting activities. It also hinders researchers from being aware of the challenges and remedies for the identified research gaps. We aim to identify, analyze, and synthesize the literature on mining architectural information in software repositories in terms of architectural information and sources mined, architecting activities supported, approaches and tools used, and challenges faced. A Systematic Mapping Study (SMS) has been conducted on the literature published between January 2006 and December 2022. Of the 104 primary studies finally selected, 7 categories of architectural information have been mined, among which architectural description is the most mined architectural information; 11 categories of sources have been leveraged for mining architectural information, among which version control system (e.g., GitHub) is the most popular source; 11 architecting activities can be supported by the mined architectural information, among which architecture understanding is the most supported activity; 95 approaches and 56 tools were proposed and employed in mining architectural information; and 4 types of challenges in mining architectural information were identified. This SMS provides researchers with promising future directions and help practitioners be aware of what approaches and tools can be used to mine what architectural information from what sources to support various architecting activities.
@@ 2 Research Context
To clarify the scope of this SMS, two fundamental concepts need to be explained, “mining software repositories” and “architectural information”.
@@ 2.1 Mining Software Repositories
The studies on mining software repositories collect and analyze the rich data available in software repositories to uncover interesting and actionable information about software projects (Hassan, 2008). VCS (e.g., GitHub), Q&A sites (e.g., Stack Overflow), and issue tracking systems (e.g., Jira), among others, are examples of sources of software repositories that are commonly used for mining development related information. These sources are useful for both software practitioners and researchers. For example, practitioners (e.g., architects) can share and learn architectural tactics and quality attributes knowledge from their peers in Stack Overflow (Bi et al., 2021). Researchers can benefit from the available data (e.g., source code, commit messages) in software repositories to conduct their research on software development, such as code clone detection (Nafi et al., 2019) and dependency analysis (Bavota et al., 2014). To achieve this, researchers are required to select software repositories and data sources that fit their research needs, extract data from these repositories, and analyze the data to obtain evidence for answering their research questions. In this SMS, we gather the literature on mining architectural information from software repositories to support the development. Specifically, we search the studies that develop approaches and tools for mining architectural information, aiming to support architecting activities (e.g., mining source code to support architecture maintenance and evolution (Kazman et al., 2015)).
@@ 2.2 Architectural Information
Architectural information states a high-level abstraction of software systems, and it is one of the most important types of information in software development (Bass et al., 2012). There are various types of architectural information, such as Architecturally Significant Requirements (ASRs), architectural solutions, and architectural decisions that are intensively explored during the development. In this section, we introduce the concept of architectural information through these three concrete types of architectural information.
@@
To create a software architecture, one must understand the problem, find a solution (architectural solution), and evaluate the final result (Souza et al., 2019). Understanding the problem requires identifying, analyzing, and prioritizing the stakeholders’ needs, producing the so-called architecturally significant requirements, such as functional and non-functional requirements (Souza et al., 2019). Architectural solutions such as architectural patterns (e.g., Model–View–Controller), tactics (e.g., resource pooling), and frameworks (e.g., Django) are the fundamental building blocks in modern software design and they are used to address the ASRs (Bass et al., 2012). Finding architectural solutions is related to deriving a software architecture from the requirements (Souza et al., 2019). However, finding a suitable architectural solution from the existing alternative solutions requires a series of design reasoning to satisfy the stakeholders’ needs and the system’s requirements (e.g., performance, security, scalability). Finally, evaluating architectural solutions requires deciding whether, and to what extent, the chosen solutions solve the problems. Architectural decision is a description of a set of architectural additions, subtractions, and modifications to the architecture, the rationale, the design rules, design constraints, and additional requirements that (partially) realize one or more requirements on a given architecture (Jansen and Bosch, 2005). Architectural decisions play a crucial role in software architecture, during the design, implementation, evolution, reuse, and integration of architectures (Jansen and Bosch, 2005).
@@
Architectural information is often described in different formats, such as textual and graphical representation (Malavolta et al., 2013), and this information is recorded in various artifacts, such as code (Granchelli et al., 2017), architecture documents (Ding et al., 2014), that are contained in diverse sources of software repositories. However, it is very challenging for architects and developers to mine and reuse architectural information that is contained in those artifacts due to, for instance, unstructured representation of this information in texts or graphs. As mentioned in Section 1, mining repositories for SA has been the subject of the architecture research community in recent years (Soliman et al., 2021b). Researchers and practitioners have developed many approaches and tools to facilitate the mining of architectural information from various artifacts that are contained in software repositories in order to support the development.
@@ 3.1 Goal and Research Questions
RQ1. What architectural information is mined in software development?
@@
SA researchers have mined different types of architectural information, such as architectural description (Chaabane et al., 2017) and system requirements (e.g., quality attributes (Bi et al., 2021)). This RQ intends to explore the types of mined architectural information to support software development. The answer of this RQ can (i) help practitioners know and utilize the types of architectural information that have been mined to support the development and (ii) help researchers be aware of the types of architectural information that have and have not been mined so that they can explore how to use the mined architectural information and/or mine the architectural information that is not yet mined.
RQ2. What sources are used to mine architectural information?
@@
Architectural information is scattered in various sources of software repositories, such as Q&A sites (e.g., Stack Overflow) (Soliman et al., 2016), issue tracking systems (e.g., Jira) (Bhat et al., 2017), and technical blogs and tutorials (Soliman et al., 2021c). Moreover, Soliman et al. (2021c) reported that there is no single source of architectural information which contains specific architectural information concepts, and certain architectural information concepts come more often in certain architectural information sources. For example, developer communities (as a prominent example of architectural information source) contain predominantly general architectural information concepts (Tang et al., 2010), such as the benefits and drawbacks of architectural solutions (Soliman et al., 2017). In contrast, issue tracking systems contain mainly architectural issues and design decisions of existing systems (Bhat et al., 2017). On the other hand, architectural solutions and rationales (e.g., benefits and drawbacks of architectural solutions) are less present within source code repositories, such as GitHub (Soliman et al., 2021c). Therefore, the selection of a source is determined by the type of information to be mined as well as the purpose of mining architectural information. For instance, information related to architectural issues or debts may require selecting issue tracking systems (e.g., Jira). Therefore, researchers have mined various sources for the sake of certain architectural information to support the development. The answer of this RQ provides insights into specific sources researchers utilize when mining architectural information, and such insights can (i) motivate other researchers to further investigate architectural information in those sources and (ii) help practitioners know where they can search architectural information to address their architectural design concerns.
RQ3. What architecting activities can be supported by the mined architectural information?
@@
There are various architecting activities (e.g., architecture understanding, architecture description, architectural implementation, architecture maintenance, architecture recovery (Li et al., 2013; Weinreich and Buchgeher, 2012)) that are performed during the architecture life cycle. Each architecting activity may require specific architectural information to be conducted effectively. The answer of this RQ can help practitioners know which architecting activities can be supported by the mined architectural information. For example, during architecture evaluation, architectural information, such as the benefits and drawbacks of certain architectural solutions (e.g., architectural patterns and tactics) can help architects evaluate the architectural solutions (e.g., choosing suitable architectural patterns according to the positively and negatively affected quality attributes) in specific application domains.
RQ4. What approaches and tools are used to mine architectural information?
@@
There might be approaches and tools that have been proposed and used to mine architectural information. The answer of this RQ can provide practitioners with an overview of the readily available approaches and tools that they can utilize to mine architectural information from software repositories, and to some extent inspire researchers to develop new and dedicated architectural information mining approaches and tools.
RQ5. What are the challenges in mining architectural information?
@@
Mining architectural information in software repositories has many challenges. With this RQ, we want to identify and report these challenges as potential future research directions in this area.
@@ 3.2.1 Phase 1: Automatic Search
ACM Digital Library
https://dl.acm.org/
Study title, abstract
IEEE Explore
https://ieeexplore.ieee.org/
Study title, abstract, keywords
Science Direct
https://www.sciencedirect.com/
Study title, abstract, keywords
EI Compendex
https://www.engineeringvillage.com/
Study title, abstract
Springer Link
https://link.springer.com/
Study title, abstract
Wiley InterScience
https://onlinelibrary.wiley.com/
Study title, abstract
ISI Web of Science
https://login.webofknowledge.com/
Study title, abstract, keywords
Scopus
https://scopus.com/
Study title, abstract, keywords
@@ 3.2.2 Phase 2: Study Screening
E4. A short paper is excluded.
@@ 3.2.3 Phase 3: Snowballing
In order to minimize the threat of missing relevant studies that were possibly missed out during the automatic search, we conducted snowballing by following the approach proposed by Wohlin (2014). Specifically, we adopted the forward (i.e., collecting those studies citing the selected studies) and backward (i.e., using the references of the selected studies) snowballing. We conducted the snowballing process iteratively, and the iterative process was completed when there were no newly selected studies. Specifically, we performed two iterations during the snowballing process. We used the set of studies selected from the final round of study screening in the previous section (see Section 3.2.2) as the start seed (i.e., 74 studies). We then looked for citations from each of the 74 studies in the start seed. We used Google Scholar to collect citations for each paper. We downloaded the citations and analyzed each one according to the inclusion and exclusion criteria defined in Table 3. In the first iteration, 3,228 citations (of the 74 studies) were found. These 3,212 citations were further screened based on the titles (248 studies retained), abstracts (53 studies retained), and full texts (27 studies retained). In the second iteration, we considered a start seed of 27 studies retained from the first iteration and 91 citations were collected. Akin to the citations obtained from the first iteration, the 91 citations were also subjected to the inclusion and exclusion criteria, and 21 studies were retained based on titles, 6 studies were identified based on abstracts, and 3 studies were finally selected based on full texts. Finally, 30 (i.e., 27+3) primary studies were added through the snowballing process. In Figure 2, we show the snowballing process that was performed in this SMS. Therefore, we have selected 104 (i.e., 74+30) primary studies in our SMS. The details of the selected 104 primary studies are provided in Appendix A.
@@ 3.2.4 Phase 4: Data Extraction
To provide the demographic information of the selected studies and answer the five RQs of this SMS (see Table 1), we carefully read all the primary studies selected in this SMS (see Figure 1) and extracted the required data items as listed in Table 4. Before the formal data extraction, the first three authors discussed the meaning of each data item and the way about how to extract data. To ensure an unambiguous understanding of the data items, the first author conducted a pilot data extraction. Specifically, the first author randomly selected 15 primary studies (from 104 selected studies, see Figure 2) and extracted the data according to the data items in Table 4. To suppress the effect of subjective bias, the pilot data extraction results were reviewed and validated by other two authors (i.e., the second and third authors) of this study so that three authors could get a consensus on the understanding of the extracted data items. Likewise, in the formal data extraction process, the data extraction was performed by the first author and validated by the second and third authors, and any divergences and ambiguities of the results in the extracted data were discussed together to reach an agreement. In this way, we can ensure that the extracted data in the formal data extraction process are valid. The description of the data items and their relevant RQs are presented in Table 4. Specifically, nine data items have been defined to be extracted from the selected studies. The three data items (i.e., D1-D3) are used to extract the demographic details of the selected studies, and the remaining data items (i.e., D4-D9) are used to answer the five RQs (see Table 1). The data extraction was subsequently followed by data synthesis, and these two processes were conducted and recorded with the aid of MAXQDA (a qualitative data analysis tool) [^1] https://www.maxqda.com/.
@@
Data item Description Data analysis approach Relevant RQ
D1
Publication year
The publication year of the study.
Descriptive statistics
Overview
D2
Publication venue
The name of the venue where the study is published.
Descriptive statistics
Overview
D3
Publication type
The type of the study (i.e., journal, conference, or workshop).
Descriptive statistics
Overview
D4
Architectural information
The type of mined architectural information.
Open coding & constant comparison, predefined classifications, and descriptive statistics
RQ1
D5
Source
The source used to mine architectural information.
Open coding & constant comparison and descriptive statistics
RQ2
D6
Dataset
The dataset used to mine architectural information.
Open coding & constant comparison and descriptive statistics
RQ2
D7
Architecting activity
The architecting activity that a study claims to support.
Open coding & constant comparison, predefined classifications, and descriptive statistics
RQ3
D8
Approach
The approach used to mine architectural information.
Open coding & constant comparison and descriptive statistics
RQ4
D9
Tool
The tool used to mine architectural information.
Open coding & constant comparison and descriptive statistics
RQ4
D10
Challenge
The challenge faced when mining architectural information.
Open coding & constant comparison and descriptive statistics
RQ5

### v4

@@ Abstract
Mining Software Repositories (MSR) has become an essential activity in software development. Mining architectural information (e.g., architectural models) to support architecting activities, such as architecture understanding, has received significant attention in recent years. However, there is a lack of clarity on what literature on mining architectural information is available. Consequently, this may create difficulty for practitioners to understand and adopt the state-of-the-art research results, such as what approaches should be adopted to mine what architectural information in order to support architecting activities. It also hinders researchers from being aware of the challenges and remedies for the identified research gaps. We aim to identify, analyze, and synthesize the literature on mining architectural information in software repositories in terms of architectural information and sources mined, architecting activities supported, approaches and tools used, and challenges faced. A Systematic Mapping Study (SMS) has been conducted on the literature published between January 2006 and December 2022. Of the 104 primary studies finally selected, 7 categories of architectural information have been mined, among which architectural description is the most mined architectural information; 11 categories of sources have been leveraged for mining architectural information, among which version control system (e.g., GitHub) is the most popular source; 11 architecting activities can be supported by the mined architectural information, among which architecture understanding is the most supported activity; 95 approaches and 56 tools were proposed and employed in mining architectural information; and 4 types of challenges in mining architectural information were identified. This SMS provides researchers with promising future directions and help practitioners be aware of what approaches and tools can be used to mine what architectural information from what sources to support various architecting activities.
@@ 2 Research Context
To clarify the scope of this SMS, two fundamental concepts need to be explained, mining software repositories and architectural information.
@@ 2.1 Mining Software Repositories
The studies on mining software repositories collect and analyze the rich data available in software repositories to uncover interesting and actionable information about software projects (Hassan, 2008). VCS (e.g., GitHub), Q&A sites (e.g., Stack Overflow), and issue tracking systems (e.g., Jira), among others, are examples of sources of software repositories that are commonly used for mining development related information. These sources are useful for both software practitioners and researchers. For example, practitioners (e.g., architects) can share and learn architectural tactics and quality attributes knowledge from their peers in Stack Overflow (Bi et al., 2021). Researchers can benefit from the available data (e.g., source code, commit messages) in software repositories to conduct their research on software development, such as code clone detection (Nafi et al., 2019) and dependency analysis (Bavota et al., 2014). To achieve this, researchers are required to select software repositories and data sources that fit their research needs, extract data from these repositories, and analyze the data to obtain evidence for answering their research questions. In this SMS, we gather the literature on mining architectural information from software repositories to support the development. Specifically, we search the studies that develop approaches and tools for mining architectural information, aiming to support architecting activities (e.g., mining code to support architecture maintenance and evolution (Kazman et al., 2015)).
@@ 2.2 Architectural Information
Architectural information states a high-level abstraction of software systems, and it is an important type of information in software development (Bass et al., 2012). There are various types of architectural information, such as Architecturally Significant Requirements (ASRs), architectural solutions, and architectural decisions that are intensively explored during the development. In this section, we introduce the concept of architectural information through these three concrete types of architectural information.
@@
To create an architecture, one must understand the problem, find an architectural solution, and evaluate the final result (Souza et al., 2019). Understanding the problem requires identifying, analyzing, and prioritizing the stakeholders’ needs, producing the so-called architecturally significant requirements, such as functional and non-functional requirements (Souza et al., 2019). Architectural solutions such as architectural patterns (e.g., Model–View–Controller), tactics (e.g., resource pooling), and frameworks (e.g., Django) are the fundamental building blocks in modern software design and they are used to address the ASRs (Bass et al., 2012). Finding architectural solutions is related to deriving a software architecture from the requirements (Souza et al., 2019). However, finding a suitable architectural solution from the existing alternative solutions requires a series of design reasoning to satisfy the stakeholders’ needs and the system’s requirements (e.g., performance, security, scalability). Finally, evaluating architectural solutions requires deciding whether, and to what extent, the chosen solutions solve the problems. Architectural decision is a description of a set of architectural additions, subtractions, and modifications to the architecture, the rationale, design rules, design constraints, and additional requirements that (partially) realize one or more requirements on a given architecture (Jansen and Bosch, 2005). Architectural decisions play a crucial role in software architecture, during the design, implementation, evolution, reuse, and integration of architectures (Jansen and Bosch, 2005).
@@
Architectural information is often described in different formats, such as textual and graphical representation (Malavolta et al., 2013), and this information is recorded in various artifacts, such as code, architecture documents (Ding et al., 2014), that are contained in diverse sources of software repositories. However, it is very challenging for architects and developers to mine and reuse architectural information that is contained in those artifacts due to, for instance, unstructured representation of this information in texts or graphs. As mentioned in Section 1, mining repositories for SA has been the subject of the architecture research community in recent years (Soliman et al., 2021b). Researchers and practitioners have developed many approaches and tools to facilitate the mining of architectural information from various artifacts that are contained in software repositories in order to support the development.
@@ 3.1 Goal and Research Questions
RQ1. What architectural information is mined in software development? SA researchers have mined different types of architectural information, such as architectural description (Chaabane et al., 2017) and system requirements (e.g., quality attributes (Bi et al., 2021)). This RQ intends to explore the types of mined architectural information to support software development. The answer of this RQ can (i) help practitioners know and utilize the types of architectural information that have been mined to support the development and (ii) help researchers be aware of the types of architectural information that have and have not been mined so that they can explore how to use the mined architectural information and/or mine the architectural information that is not yet mined.
@@
RQ2. What sources are used to mine architectural information? Architectural information is scattered in various sources of software repositories, such as Q&A sites (Soliman et al., 2016), issue tracking systems (Bhat et al., 2017), and technical blogs and tutorials (de Dieu et al., 2022). There is no single source of architectural information that contains all architectural information concepts, and certain architectural information concepts come more often from specific architectural information sources (Soliman et al., 2021c). For example, developer communities (as an example of architectural information source) contain predominantly general architectural information concepts, such as the benefits and drawbacks of architectural solutions (Soliman et al., 2017). In contrast, issue tracking systems contain mainly architectural issues and design decisions of existing systems (Bhat et al., 2017). On the other hand, architectural solutions and rationales (e.g., benefits and drawbacks of architectural solutions) are less present within source code repositories, such as GitHub (Soliman et al., 2021c). Therefore, researchers have mined various sources for the sake of certain architectural information to support the development. The answer of this RQ provides insights into specific sources researchers utilize when mining architectural information, and such insights can (i) motivate other researchers to further investigate architectural information in those sources and (ii) help practitioners know where they can search architectural information to address their architectural design concerns.
@@
RQ3. What architecting activities can be supported by the mined architectural information? There are various architecting activities (e.g., architecture understanding, architecture description, architectural implementation, architecture maintenance, architecture recovery (Li et al., 2013; Weinreich and Buchgeher, 2012)) that are performed during the architecture life cycle. Each architecting activity may require specific architectural information to be conducted effectively. The answer of this RQ can help practitioners know which architecting activities can be supported by the mined architectural information. For example, during architecture evaluation, architectural information, such as the benefits and drawbacks of certain architectural solutions (e.g., architectural patterns and tactics) can help architects evaluate the architectural solutions (e.g., choosing suitable architectural patterns according to the positively and negatively affected quality attributes) in specific application domains.
@@
RQ4. What approaches and tools are used to mine architectural information? There might be approaches and tools that have been proposed and used to mine architectural information. The answer of this RQ can provide practitioners with an overview of the readily available approaches and tools that they can utilize to mine architectural information from software repositories, and to some extent inspire researchers to develop new and dedicated architectural information mining approaches and tools.
@@
RQ5. What are the challenges in mining architectural information? Mining architectural information in software repositories has many challenges, such as vagueness or ambiguity in architectural element description and applicability of approaches to heterogeneous architectural datasets. With this RQ, we want to identify and report these challenges as potential future research directions in this area.
@@ 3.2.1 Phase 1: Automatic Search
ACM Digital Library https://dl.acm.org/ Study title, abstract
IEEE Explore https://ieeexplore.ieee.org/ Study title, abstract, keywords
Science Direct https://www.sciencedirect.com/ Study title, abstract, keywords
EI Compendex https://www.engineeringvillage.com/ Study title, abstract
Springer Link https://link.springer.com/ Study title, abstract
Wiley InterScience https://onlinelibrary.wiley.com/ Study title, abstract
ISI Web of Science https://login.webofknowledge.com/ Study title, abstract, keywords
Scopus https://scopus.com/ Study title, abstract, keywords
@@ 3.2.2 Phase 2: Study Screening
E4. A short paper (i.e., less than 4 pages) is excluded.
@@ 3.2.3 Phase 3: Snowballing
In order to minimize the threat of missing relevant studies that were possibly missed out during the automatic search, we conducted snowballing by following the approach proposed by Wohlin (2014). Specifically, we adopted the forward (i.e., collecting those studies citing the selected studies) and backward (i.e., using the references of the selected studies) snowballing. We conducted the snowballing process iteratively, and the iterative process was completed when there were no newly selected studies. Specifically, we performed two iterations during the snowballing process. We used the set of studies selected from the final round of study screening in the previous section (see Section 3.2.2) as the start seed (i.e., 74 studies). We then looked for citations from each of the 74 studies in the start seed. We used Google Scholar to collect citations for each paper. We downloaded the citations and analyzed each one according to the inclusion and exclusion criteria defined in Table 3. In the first iteration, 3,228 citations (of the 74 studies) were found. These 3,212 citations were further screened based on the titles (248 studies retained), abstracts (53 studies retained), and full texts (27 studies retained). In the second iteration, we considered a start seed of 27 studies retained from the first iteration and 91 citations were collected. Akin to the citations obtained from the first iteration, the 91 citations were also subjected to the inclusion and exclusion criteria, and 21 studies were retained based on titles, 6 studies were identified based on abstracts, and 3 studies were finally selected based on full texts. Finally, 30 (i.e., 27+3) primary studies were added through the snowballing process. In Figure 2, we show the snowballing process that was performed in this SMS. Therefore, we have selected 104 (i.e., 74+30) primary studies in our SMS. The details of the selected 104 primary studies are provided in the Selected Studies file in the Supplementary Material (de Dieu et al., 2024).
@@ 3.2.4 Phase 4: Data Extraction
To provide the demographic information of the selected studies and answer the five RQs of this SMS (see Table 1), we carefully read all the primary studies selected in this SMS (see Figure 1) and extracted the required data items as listed in Table 4. Before the formal data extraction, the first three authors discussed the meaning of each data item and the way about how to extract data. To ensure an unambiguous understanding of the data items, the first author conducted a pilot data extraction. Specifically, the first author randomly selected 15 primary studies (from 104 selected studies, see Figure 2) and extracted the data according to the data items in Table 4. To suppress the effect of subjective bias, the pilot data extraction results were reviewed and validated by other two authors (i.e., the second and third authors) of this study so that three authors could get a consensus on the understanding of the extracted data items. Likewise, in the formal data extraction process, the data extraction was performed by the first author and validated by the second and third authors, and any divergences and ambiguities of the results in the extracted data were discussed together to reach an agreement. In this way, we can ensure that the extracted data in the formal data extraction process are valid. The description of the data items and their relevant RQs are presented in Table 4. Specifically, nine data items have been defined to be extracted from the selected studies. The three data items (i.e., D1-D3) are used to extract the demographic details of the selected studies, and the remaining data items (i.e., D4-D9) are used to answer the five RQs (see Table 1). The data extraction was subsequently followed by data synthesis, and these two processes were conducted and recorded with the aid of MAXQDA (a qualitative data analysis tool).
@@
# Data item Description Data analysis approach Relevant RQ
D1 Publication year The publication year of the study. Descriptive statistics Overview
D2 Publication venue The name of the venue where the study is published. Descriptive statistics Overview
D3 Publication type The type of the study (i.e., journal, conference, or workshop). Descriptive statistics Overview
D4 Architectural information The type of mined architectural information. Open coding & constant comparison, predefined classifications, and descriptive statistics RQ1
D5 Source The source used to mine architectural information. Open coding & constant comparison and descriptive statistics RQ2
D6 Dataset The dataset used to mine architectural information. Open coding & constant comparison and descriptive statistics RQ2
D7 Architecting activity The architecting activity that a study claims to support. Open coding & constant comparison, predefined classifications, and descriptive statistics RQ3
D8 Approach The approach used to mine architectural information. Open coding & constant comparison and descriptive statistics RQ4
D9 Tool The tool used to mine architectural information. Open coding & constant comparison and descriptive statistics RQ4
D10 Challenge The challenge faced when mining architectural information. Open coding & constant comparison and descriptive statistics RQ5